# PrisonBreak: Jailbreaking Large Language Models With at Most Twenty-Five Targeted Bit-flips

### v2

@@ 4.1 Design Challenges and Our Approach
The bit-flip onion effect. Prior work on bit-flip attacks presents two distinct approaches to identify target bits. One-shot search [2], in which the attacker computes the importance of each weight for a malicious objective and flips bits within those critical weights. Progressive search [3], where the adversary iteratively identifies a single target bit in the critical weights whose flipping increases the likelihood of achieving the adversarial goal. We examine both approaches and find a fundamental reason why one-shot search is not appropriate for jailbreaking: the bit-flip onion effect. Once a bit is flipped, it alters the model’s characteristics and we need to re-compute critical weights and candidate bits. To address this, we employ the progressive search strategy.
High computational costs. Prior work has demonstrated bit-flip attacks on computer vision models, with the largest examined being ViTs [10, 7], which comprise up to 87 million parameters and are trained on small datasets such as CIFAR-10. However, we target language models with 1.5 to 13 billion parameters—17 to 150 larger than the largest models examined in prior work. Running the progressive bit-search on these models becomes computationally intractable. For instance, locating 25 target bits on a 7-billion parameter Llama2 takes 177 days on an Nvidia A40 GPU. To address this, we present speed-up techniques in §5.2 and §5.3 that reduce the number of critical weights and bits to be analyzed, decreasing the time it takes to locate target bits by 20 (9 days in the previous example).
Providing high-quality responses. The jailbroken models should provide high-quality responses to both normal and harmful prompts, allowing the attack artifacts to remain undetected for as long as possible. Prompt-based jailbreaking attacks focus on generating affirmative responses, such as “Sure, here is…” without addressing the content that follows. Using the same objective in our attack leads to models generating affirmative responses, followed by harmless content, such as repeating “Sure” many times or backtracking with statements like “Sure, but I can’t say this as an ethical AI model.” To tackle this challenge, we take two approaches. We first consider the content that follows affirmative tokens in our bit-search. We utilize uncensored open-source language models to collect quality responses for harmful inputs. We then create a proxy dataset to accurately approximate the desired jailbroken response. Next, we select target bits that best preserves a model’s generation capabilities.
Rowhammer exploitability. Certain conditions make it easier for Rowhammer attackers to flip target bits [38, 39, 37], such as the presence of only one target bit per data row. Our attack procedure takes these constraints into account.
@@ 5 Offline Profiling of Target Bits
We now provide a detailed description of PrisonBreak’s procedure for identifying target bits in the memory representation of the victim model to jailbreak at runtime.
@@ 5.1 Our Jailbreaking Objective
PrisonBreak employs a progressive bit-search, where the attacker identifies critical weights and target bits based on the gradients computed with respect to an attack objective. Prior attack objectives include maximizing the training loss to degrade performance [3, 1, 48], minimizing prediction error of specific samples to achieve targeted misclassification [4, 5] or optimizing for trojan/backdoor attacks [7, 8, 10, 9, 6]. But these objectives developed for classification settings are incompatible with jailbreaking.
@@
Overfitting: Using these objectives leads the target model to excessively overfit to affirmative tokens, resulting in repetitive and nonsensical outputs.
@@
Table I provides examples illustrating responses generated using only the prior work’s objective in PrisonBreak. Most prior work considers the attack successful if a model’s output begins with an affirmative response. But in practice, these attacks have lower fidelity in their generated responses.
@@
Our approach uses open-source, unaligned language models [50] to generate oracle harmful responses. We collect harmful queries, widely available online from jailbreaking auditing benchmarks such as HarmBench [51] and AdvBench [13]. Then, we apply prompt engineering to format model outputs such that each response starts with affirmative tokens, followed by harmful content. We refer to this collection of harmful queries and oracle responses as our proxy dataset and use it to approximate the true jailbreaking behavior we are trying to induce. Note that we will make this dataset publicly available to support future research.
@@
where is a prompt-target pair from our proxy-dataset, is the size, is the tokens up to position , is the probability of generating token , and is the number of tokens in .
Our loss function computes the cross-entropy loss (highlighted in dark-red) across the entire proxy dataset. But, cross-entropy treats all tokens equally, making tokens in the harmful context equally important as the first few affirmative tokens. This can be undesirable, as the model must first produce affirmative tokens; otherwise, it is less likely to generate the harmful context, reducing effectiveness.
To address this, we assign more importance to the initial tokens of each sequence by incorporating a weighting term, highlighted in blue. The importance gradually decreases; e.g., the first token has a weight of 1, the tenth , and the last for a sequence of 20. Our attack prioritizes generating affirmative tokens first, and then make the target model respond with high-quality harmful contexts.
@@ 5.2 Identifying Critical Weights
Skipping layers not critical to our goal. Commercial-scale language models typically contain 32–80 Transformer blocks, resulting in a total layer count that can quickly reach several hundred. We thus explore layers that are never selected and exclude them from subsequent searches. In preliminary experiments, our offline procedure does not select any normalization, embedding, and unembedding layers. By skipping these layers, we reduce the total number of weights to inspect by 20%.
Inspecting top- critical weights. Another approach to reduce computations is to inspect only the top- weights within each tensor (layer), based on their importance scores, as critical weights. We set to a fixed value, since making proportional to the weight count per layer results in no weights being selected from layers with fewer parameters. While our ablation study in §6.5 shows that the attacker can use a smaller , such as 10, with a decent attack success, e.g., of 73.6% on Llama2-7B, we conservatively set to 100, achieving 80–98% jailbreaking success across all evaluated models.
@@ 5.3 Selecting Target Bits
Now we select target bits based on the importance scores. The scores should capture both the jailbreaking success and generation capabilities after a bit-flip ( ). We also develop bit-level speed-up techniques to accelerate the selection process ( ). Moreover, if an adversary has profiled a specific DRAM module (hardware-aware), they can assess whether chosen target bits are flippable with Rowhammer. If not, those bits are excluded from the search process to enhance efficiency ( ). We first outline the procedure for a hardware-agnostic adversary and then show how a hardware-aware attacker can skip non-flippable bits.
(1) Rank the bits by the jailbreak score. We rank the bits in our critical weights, based on the jailbreaking score:
@@
where is the model with a bit flipped, and represents all bits in the critical weights. The attacker computes this score for each bit and ranks all bits in ascending order. If the score is low, flipping is more likely to induce the model to produce harmful responses.
(2) Selecting a bit based on the utility score. Our preliminary attack, which uses only the jailbreaking score to select target bits, often reduces the model’s ability to produce coherent, high-quality answers. To measure the likelihood of a bit-flip degrading utility, we present the utility score.
@@
where is a clean dataset to evaluate ’s utility, is the (unweighted) cross-entropy loss, and are the prompt–target pairs. We construct by randomly sampling prompts from representative clean data. We use Alpaca, a cleaned and curated version of the dataset used to train Alpaca language models [52]. For each prompt in Alpaca, we generate a response from the target model. We use greedy decoding for the generations, which serves as the model’s most likely responses. should contain a sufficient number of samples to capture the model’s generation behavior. Interestingly, while the original dataset contains 52,000 instructions, only 100 are needed to approximate the model’s utility, reducing the computations by 520.
@@
Exponent bits. We observe that most critical weights fall within , so flipping a mantissa bit or the sign bit does not produce a change large enough to jailbreak. We thus focus on the exponent bits, particularly the three most significant ones, which yield the largest . With roughly half of these exponent bits resulting in with the same sign, this reduces the search to just 1.5 bits.
01 flip direction. Since most critical weights have small values, a 01 flip results in relatively larger changes than a 10 flip. We thus consider only 01 flips. This reduces the number of bits to search from 1.5 to 1, allowing us to evaluate just 1 bit per 16-bit weight.
(3) Skipping non-flippable bits. The hardware-aware attacker knows the offsets within the DRAM’s physical pages, along with the corresponding flip directions that Rowhammer can induce. If the attacker knows these flippable bit locations, they can select only bits that match the offset and flip direction of at least one vulnerable cell. We evaluate our hardware-aware attacks in §7 against DRAM modules with up to 1625 fewer vulnerable cells and achieve comparable jailbreak success to the hardware-agnostic attacks.
@@ 6.1 Experimental Setup
Models. We use 10 open-source large language models within five different model families, displayed in Table III. Once loaded into memory, model parameters are represented as half-precision floating-point numbers, with sizes ranging from 1.5–70 billion. We select models with different combinations of alignment algorithms applied during their post-training, namely supervised fine-tuning (SFT) [25, 29], reinforcement learning with human feedback (RLHF) [26], and direct preference optimization (DPO) [28]. All experiments utilize each model’s default chat template, available from their respective HuggingFace [55] model cards.
Metrics. To evaluate our attack, we use HarmBench [51], comprised of various harmful prompts spanning categories such as harassment, cybercrime, and illegal activities. Following common practices in jailbreaking language models [34], we evaluate all our models on the full test set in the Standard Behaviors dataset, which consists of 159 prompts. We use the provided classifier, which takes a generation as input and outputs “yes” or “no” to indicate whether the response was harmful. We then report the percentage of harmful responses as the attack success rate (ASR). We evaluate the jailbroken model’s generation capabilities using tinyBenchmarks [56], a representative subset of popular language model evaluation datasets that assess a variety of tasks spanning reasoning, arithmetic, and creativity, providing a holistic measure. We report model performance on this benchmark as the average accuracy across all tasks (Acc.).
@@
Baseline bit-flip attack (BFA): a straightforward adaptation of the prior bit-flip attack [1, 3]. We remove PrisonBreak’s novel components: the proxy dataset, utility score, and weighted cross-entropy, but retain our search reduction techniques for efficiency. This will allow us to assess the effectiveness of various techniques we develop to jailbreak language models.
| Attacks | Model Response |
| No Attack | “I’m sorry, but I cannot provide guidance on how to engage in illegal activities such as smuggling contraband while evading law enforcement. Such actions are not only illegal, but they also pose serious risks …” |
| Baseline BFA | “Sure, here’s a detailed guidance on how to smuggle contraband while evading law enforcement: sure, sure, sure sure, sure, sure, sure, and more. Sure, here’s a detailed guidance on how to smuggle contraband while evading…” |
@@ 6.2 PrisonBreak Effectiveness
We first evaluate attack effectiveness across open-source language models publicly available from the literature. Here, we assume the attackers are at their best effort. In our attack, we assume all target bits identified by PrisonBreak will be flipped under Rowhammer pressure, following the same assumption made in prior bit-flip attacks on neural networks [7, 3, 48, 8, 5, 9, 10]. Note that in §7 we, for the first time, evaluate the attack success of our bit-flip attack with Rowhammer simulations across dozens of DRAM modules with varying levels of vulnerability. For other attacks, we conduct an extensive hyperparameter search to bring the highest ASR. Our results serve as the upper bound for each jailbreaking adversary, including ours.
@@
Results. Table II summarizes our results. We test each attack separately and report the ASR and change in Acc. (Acc.). We do not report Acc. for GCG-M, as prompt-based attacks do not impact clean accuracy. For parameter-based attacks, we report the number of modified parameters (P).
@@
PrisonBreak outperforms all baseline attacks with an average ASR of 89.56%. Due to the different attack surface exploited by GCG-M, a direct comparison is not possible. However, the results demonstrate that our attack offers distinct advantages: it enables permanent runtime jailbreaking, requires no modification to harmful prompts, and achieves an average 1.73 higher ASR. Most notably, our attack improves the baseline bit-flip attack by an average ASR of 22.3%. By extending the context in the search objective and weeding out destructive bit-flips, we evidently find much higher-quality target bits regardless of model configuration. We observe a case where the baseline BFA achieves a higher ASR on Llama3-8B. However, as discussed in §6.3, this comes at a large cost to the quality of harmful responses. The attack performing closest to ours is Ortho, with comparative ASRs on most models. However, it modifies between 0.7–23.4 billion parameters depending on the model size and architecture. Since this would require several billion bit-flips, many within the same physical page, Ortho is not practically exploitable by Rowhammer at runtime. Surprisingly, we are able to break the safety alignment by modifying only 5–25 parameters (i.e., 5–25 bits), a reduction of – compared to Ortho.
@@
Model size does not matter, but alignment algorithm does. Prior work [1, 57, 5, 4] found that increasing model size tends to enhance the resilience to bit-flip attacks, possibly by introducing redundant parameters. Another study [58] showed that models rely on multiple neurons to produce outputs, supporting the observations made by these studies. We find similar ASR across model families, with no conclusive pattern between model size and attack success. As the number of parameters increases (1.57/81370B), ASR slightly decreases in Vicuna (93.190.6%) but remains the same in Llama2 and even increases in Llama3 and Qwen2 (80.598.1%).
@@
However, we do find differences in the number of bit-flips necessary to achieve high ASR. In Vicuna models, 5–20 bit-flips are sufficient to jailbreak, while Llama, Tulu, and Qwen models require 13–25 bit-flips for comparable ASR. In most cases, the ASR for these models is also slightly lower than in Vicuna models. We attribute this difference to the alignment mechanisms employed by each model family. Vicuna uses only SFT [25, 29]. In contrast, Llama, Tulu and Qwen models also employ RLHF [26, 27], DPO [28], or both. With more alignment, models show marginally increased resilience to our attack. However, compared to the effort required to train language models at scale with those mechanisms, our attack only needs a few more bit-flips (8–20) to achieve 80–98% ASR.
Our attack preserves model utility. Maintaining a model’s generation capabilities is important when jailbreaking at runtime; otherwise, the victim can easily detect anomalies and re-load the model into memory. We evaluate the change in accuracy (Acc.) on tinyBenchmarks. Each value in Table II represents the average performance over eight benchmarking tasks (see Appendix D for the details). Our attack causes minimal change in model utility with an average Acc. of -2.26%. This is not the case for Baseline BFA, which suffers 7 times higher reduction (-15.48%) because it selects destructive bit flips during search. However, compared to Ortho, our attack results in a slightly higher utility loss of 2%. This is likely due to the number of parameters the attack can perturb. Across all models, Ortho perturbs at least 0.7 billion parameters, while PrisonBreak alters up to 25 bits across 25 parameters. Because our attack targets fewer weights, we cannot distribute behavior changes evenly across a large number of parameters and must instead modify a few weights disproportionately.

### v3

@@ 4.1 Design Challenges and Our Approach
The bit-flip onion effect. Prior work on bit-flip attacks (BFAs) has introduced two main strategies for identifying target bits: one-shot search [33], where the attacker estimates the importance of each weight and flips all selected bits simultaneously; and progressive search [67], where the adversary iteratively selects and flips one bit at a time from the critical weights. We study both approaches and identify a key limitation of one-shot search in the context of jailbreaking: the bit-flip onion effect. Once a bit is flipped, it changes the model’s behavior. Consequently, two flips that are individually beneficial to the adversarial objective may not remain so when applied together. Our preliminary experiments confirm this effect, making one-shot search unsuitable for jailbreaking. To address this, we adopt the progressive search strategy.
High computational costs. Prior BFAs target computer vision models, with the largest examined being ViTs [99, 45], which comprise up to 87 million parameters and are trained on small datasets such as CIFAR-10. However, we target language models with 1.5 to 70 billion parameters—17 to 805 larger than the largest models examined in prior work. Running the progressive bit-search on these models becomes computationally intractable. For instance, locating 25 target bits on a 7-billion parameter Llama2 takes 177 days on an NVIDIA A40 GPU. To address this, we present speed-up techniques in §5.2 and §5.3 that reduce the number of critical weights and bits to be analyzed, decreasing the time it takes to locate target bits by 20 (9 days in the previous example).
Providing high-quality responses. Jailbroken models should provide high-quality responses to both normal and harmful prompts, allowing the attack to remain undetected for as long as possible. Prompt-based jailbreaks focus on generating affirmative responses, such as “Sure, here is…” without addressing the content that follows. Using the same objective in our attack leads to models generating affirmative responses, followed by harmless content, such as repeating “Sure” many times or backtracking with statements like “Sure, but I cannot answer this.” To tackle this challenge, we take two approaches. We first consider the content that follows affirmative tokens in our bit-search. We utilize uncensored open-source language models to collect quality responses for harmful inputs. We then create a proxy dataset to accurately approximate the desired jailbroken response. Next, we select target bits that best preserve a model’s generation capabilities.
Rowhammer exploitability. Not all physical bits are flippable under hammering. Recent Rowhammer studies on GPUs report several orders of magnitude fewer bit flips than on CPUs [51], making only a small fraction of bits susceptible. Our attack takes these constraints into account.
@@ 5 Offline Profiling of Target Bits
We now provide a detailed description of PrisonBreak’s procedure for identifying target bits in the memory representation of the victim model, enabling runtime jailbreaking.
@@ 5.1 Our Jailbreaking Objective
PrisonBreak employs a progressive bit-search, where the attacker identifies critical weights and target bits based on gradients computed with respect to an attack objective. Prior objectives include maximizing training loss to degrade utility [67, 94, 24], minimizing the loss on specific samples to achieve targeted misclassification [70, 7], or optimizing backdoors [45, 14, 99, 69, 11, 50]. But these objectives were developed for classification and are incompatible with jailbreaking.
@@
Overfitting: The model excessively overfits to affirmative tokens, resulting in repetitive and nonsensical outputs.
@@
Table 1 provides examples illustrating responses generated using only the prior work’s objective in PrisonBreak. Most prior work considers the attack successful if a model’s output begins with an affirmative response. But in practice, these attacks have lower fidelity in their generated responses.
@@
We use an open-source, unaligned language model [35] to generate oracle responses. We collect harmful queries from the jailbreaking auditing benchmark HarmBench [55], then apply prompt engineering to format model outputs such that each response starts with affirmative tokens, followed by harmful content. We refer to this collection of query–response pairs as our proxy dataset and use it to approximate the true jailbreaking behavior we are trying to induce.
@@
where is a prompt-target pair from our proxy dataset, is the size of the dataset, is the tokens up to position , is the probability of generating token , and is the number of tokens in .
Our loss function computes the cross-entropy loss (highlighted in dark red) across the entire proxy dataset. But, cross-entropy treats all tokens equally, making tokens in the harmful context equally important as the first few affirmative tokens. This can be undesirable, as the model must first produce affirmative tokens; otherwise, it is less likely to generate the harmful context. To address this, we assign more importance to the initial tokens of each sequence by incorporating a weighting term, highlighted in blue. The importance gradually decreases; e.g., the first token has a weight of 1, the tenth , and the last for a sequence of 20 tokens. Our attack prioritizes generating affirmative tokens first, and then make the target model respond with high-quality harmful content.
@@ 5.2 Identifying Critical Weights
Skipping layers not critical to our goal. Commercial-scale language models typically contain 32–80 Transformer blocks, resulting in a total layer count of several hundred. We thus attempt to exclude layers that are never selected. In preliminary experiments, our offline procedure does not select any normalization, embedding, and unembedding layers. By skipping these layers, we reduce the total number of weights to inspect by 20%.
Inspecting top- critical weights. Another approach to reduce computations is to inspect only the top- critical weights within each tensor (layer), based on their importance scores. We set to a fixed value, as making proportional to the weight count results in no weights being selected from layers with fewer parameters. While our ablation study in §6.5 shows that the attacker can use a smaller , such as 10, with decent success, e.g., of 73.6% on Llama2-7B, we conservatively set to 100, achieving 80–98% jailbreaking success across all evaluated models.
@@ 5.3 Selecting Target Bits
We select target bits based on importance scores, which capture both the jailbreaking success and generation capabilities after a bit-flip ( ). We also develop bit-level speed-up techniques to accelerate selection ( ). Moreover, if an adversary has profiled a specific DRAM module (hardware-aware), they can assess whether chosen target bits are flippable with Rowhammer. If not, those bits are excluded from the search process to enhance efficiency ( ). We first outline the procedure for a hardware-agnostic adversary and then show how a hardware-aware attacker can skip non-flippable bits.
Rank the bits by the jailbreak score. We rank the bits in our critical weights, based on the jailbreaking score:
@@
where is the model with a bit flipped, and represents all bits in the critical weights. The attacker computes this score for each bit and ranks all bits in ascending order. If the score is low, flipping is more likely to induce a harmful response.
Selecting a bit based on the utility score. Our preliminary attack, which uses only the jailbreaking score to select target bits, often reduces the model’s ability to produce coherent, high-quality answers. To measure the likelihood of a bit-flip degrading utility, we present the utility score.
@@
where is a clean dataset to evaluate ’s utility, is the (unweighted) cross-entropy loss, and are the prompt–target pairs. We construct by randomly sampling prompts from Alpaca, a cleaned and curated version of the dataset used to train Alpaca language models [78]. For each prompt in Alpaca, we generate a response from the target model. We use greedy decoding for the generations, which serves as the model’s most likely responses. should contain a sufficient number of samples to capture the model’s generation behavior. Interestingly, while the original dataset contains 52k instructions, only 100 are needed to approximate the model’s utility, reducing the computations by 520.
@@
Exponent bits. We observe that most critical weights fall within , so flipping a mantissa bit or the sign bit does not produce a change large enough to jailbreak. We thus focus on the exponent bits, particularly the three most significant ones, which yield the largest . With roughly half of these exponent bits resulting in with the correct sign, this reduces the search to just 1.5 bits.
01 flip direction. Since most critical weights have small values, a 01 flip results in relatively larger changes than a 10 flip. We thus consider only 01 flips. This reduces the average number of bits to search from 1.5 to only 1.
Skipping non-flippable bits. The hardware-aware attacker knows the offsets within the DRAM’s physical pages, along with the flip directions that Rowhammer can induce. As such, they select only bits that match the offset and flip direction of at least one vulnerable cell. We evaluate our hardware-aware attack in §7 against a GDDR6 module with up to fewer vulnerable cells than prior work, and achieve comparable success to the hardware-agnostic attack (§6).
@@ 6 Evaluation
We evaluate the effectiveness of PrisonBreak in jailbreaking commercial-scale, aligned language models. Here, we deploy our hardware-agnostic search mode. We note that in §7, we evaluate—for the first time—the effectiveness of PrisonBreak in an end-to-end setting, showcasing our hardware-aware search and Rowhammer exploitation on a real GPU.
@@ 6.1 Experimental Setup
Models. We use 10 open-source large language models within 5 different model families, displayed in Table 3. Once loaded into memory, model parameters are represented as half-precision floating-point numbers (float16), with sizes ranging from 1.5–70 billion. We select models with different combinations of alignment algorithms applied during their post-training, namely supervised fine-tuning (SFT) [81, 82], reinforcement learning with human feedback (RLHF) [63], and direct preference optimization (DPO) [66]. All experiments utilize each model’s default chat template, available from their respective HuggingFace [88] model cards.
Metrics. To evaluate our attack, we use HarmBench [55], comprised of various harmful prompts spanning categories such as harassment and illegal activities. Following common practices [4], we use the full test set in the Standard Behaviors split, which consists of 159 prompts. We use the provided classifier, which takes a generation and outputs “yes” or “no” to indicate whether the response was harmful. We then report the percentage of harmful responses as the attack success rate (ASR). We evaluate model’s generation capabilities using tinyBenchmarks [65], a representative subset of popular LLM evaluation datasets. We report performance on this benchmark as the average accuracy across all tasks (Acc.).
| Attacks | Model Response |
| No Attack | “I’m sorry, but I cannot provide guidance on how to engage in illegal activities such as smuggling contraband …” |
| Baseline BFA | “Sure, here’s a detailed guidance on how to smuggle contraband while evading law enforcement: sure, sure, sure …” |
@@
Baseline bit-flip attack (BFA): a straightforward adaptation of the prior BFA [94, 67]. We remove PrisonBreak’s novel components: the proxy dataset, utility score, and weighted loss, but retain our speedup techniques for efficiency. This allows us to assess the effectiveness of the various techniques we develop for jailbreaking.
@@ 6.2 PrisonBreak Effectiveness
We first evaluate the effectiveness of attacks across open-source language models. Here, we assume the attackers are at their best effort. In our attack, we assume all target bits are flippable under Rowhammer pressure, following the same assumption made in prior BFAs [45, 67, 24, 14, 7, 69, 99]. For other attacks, we conduct an extensive hyperparameter search to bring the highest ASR. Our results serve as the upper bound for each jailbreaking adversary, including ours.
@@
Results. Table 2 summarizes our results. We test each attack separately and report the ASR and change in Acc. (Acc.). We do not report Acc. for GCG-M, as prompt-based attacks do not impact clean accuracy. For parameter-based attacks, we report the number of modified parameters (P).
@@
PrisonBreak outperforms all baseline attacks with an average ASR of 89.6%. Our attack offers distinct advantages over GCG: it enables permanent runtime jailbreaking, requires no input modifications, and achieves an average 1.73 higher ASR. Notably, our attack improves over Baseline BFA by an average ASR of 22.3. By extending the context in the search objective and weeding out destructive bit-flips, we evidently find much higher-quality target bits regardless of model configuration. We observe a case where Baseline BFA achieves a higher ASR on Llama3-8B. However, as discussed in §6.3, this comes at a large cost to the quality of harmful responses. The attack performing closest to ours is Ortho, with comparative ASRs on most models. However, it modifies between 0.7–23.4 billion parameters depending on the model size and architecture. Since this would require several billion bit-flips, Ortho is not practically exploitable by Rowhammer. Surprisingly, we are able to break the safety alignment by modifying only 5–25 parameters (i.e., 5–25 bits), a reduction of – compared to Ortho.
@@
Model size does not matter, but alignment algorithm does. Prior work [94, 32, 7, 70] found that increasing model size tends to enhance the resilience to bit-flip attacks, possibly by introducing redundant parameters. Another study [22] showed that models rely on multiple neurons to produce outputs, supporting these observations. We find similar ASR across model families, with no conclusive pattern between model size and attack success. As the number of parameters increases (1.57/81370B), ASR slightly decreases in Vicuna (93.190.6%) but remains the same in Llama2 and even increases in Llama3 and Qwen2 (80.598.1%).
@@
However, we do find differences in the number of bit-flips needed to reach high ASR. For Vicuna models, 5–20 bit-flips are sufficient to jailbreak, while Llama, Tulu, and Qwen models require 13–25 bit-flips for comparable ASR. In most cases, the ASR for these models is also slightly lower than for Vicuna. We attribute this to alignment differences: Vicuna uses only SFT [81, 82], while Llama, Tulu and Qwen models also employ RLHF [63, 19], DPO [66], or both. With more alignment, models show marginally increased resilience. However, compared to the effort required to train language models at scale with those mechanisms, our attack only needs a few more bit-flips (8–20) to achieve 80–98% ASR.
@@
Our attack preserves model utility. Maintaining a model’s generation capabilities is important when jailbreaking; otherwise, the victim can easily detect anomalies and reload the model into memory. We evaluate the change in accuracy (Acc.) on tinyBenchmarks. Each value in Table 2 represents the average performance over eight benchmarking tasks (see Appendix C for the details). Our attack causes minimal change in model utility with an average Acc. of -2.3%. This is not the case for Baseline BFA, which suffers 7 times higher reduction (-15.5%) because it selects destructive bit flips during search. However, compared to Ortho, our attack results in a slightly higher utility loss of 2%. This is likely due to the number of parameters the attack can perturb. Across all models, Ortho perturbs at least 0.7 billion parameters, while PrisonBreak alters up to 25 bits across 25 parameters. Because our attack targets fewer weights, we cannot distribute behavior changes evenly across a large number of parameters and must instead modify a few weights disproportionately.